# On the Fundamental Limits of LLMs at Scale

### v1

@@ 6.3.2 Grounding and Reasoning Failures
Even when MLLMs integrate textual and other information (image, video, audio, etc.), the learned associations are predominantly spurious (Hosseini et al., 2025), which means that the associations are mere correlations in training data rather than causal or compositional structures. This manifests most clearly in spatial reasoning tasks, where models learn frequent co-occurrences (e.g., "cat on sofa") but not the underlying geometric or physical relations (Rahmanzadehgervi et al., 2025; Hou et al., 2025). Let $\mathcal{S}=\{o_{1},\ldots,o_{n},r_{ij}\}$ denote a visual scene with objects $o_{i}$ and spatial relations $r_{ij}$ (e.g., “on top of,” “left of,” “inside”). An ideally grounded model would learn a representation $\phi(\mathcal{S})$ that factorizes as:
@@
where $\otimes$ denotes compositional combination and $\phi_{\text{rel}}$ explicitly encodes geometric constraints (e.g., vertical displacement for ”on top of”). However, MLLM embeddings instead capture:
$\psi(\mathcal{S})\approx\sum_{i,j}w_{ij}\cdot\mathbb{1}[\text{co-occur}(o_{i},o_{j})]$
where $w_{ij}$ are learned weights reflecting dataset co-occurrence frequencies and $\mathbb{1}[\text{co-occur}(o_{i},o_{j})]$ indicates whether objects $o_{i}$ and $o_{j}$ appear together in training images.
@@
This spurious grounding is evident in compositional reasoning benchmarks. When presented with novel spatial configurations (e.g., "a sofa on top of a cat", inverting the typical arrangement), MLLMs produce nonsensical outputs or refuse to generate descriptions, because the training data contains no instances of this configuration (Thrush et al., 2022; Yuksekgonul et al., 2023). The model learns a statistical pattern, not a geometric relationship that can be flexibly recombined. Research on compositional generalization shows that multimodal models often perform no good when tested on compositional reasoning tasks that require understanding novel combinations of familiar concepts (Thrush et al., 2022). This limitation arises because the number of possible multi-way combinations grows exponentially, while training coverage is inherently limited (Lake et al., 2017; Bahdanau et al., 2019). It has been shown that even when models are explicitly trained using reinforcement learning (RL), compositional gaps with visual inputs remain (Li et al., 2025b). The failure is architectural: adding visual tokens simply expands the space of spurious correlations without imposing compositional structure.
@@
$\mathrm{Var}[Y\mid V,T]=\mathrm{Var}_{V}[Y\mid V]+\mathrm{Var}_{T}[Y\mid T]+2\,\mathrm{Cov}_{V,T}[Y]$
@@
Heavy reliance on synthetic captions or web-scraped image-text pairs causes MLLMs to overfit anthropomorphic and stylistic priors, propagating cognitive biases Shumailov et al. (2023). Human-annotated captions often impose subjective interpretations (e.g., “the dog looks happy”), while synthetic captions generated by earlier models amplify errors across generations. As a result, MLLMs reproduce these biases, producing homogenized outputs that reflect training artifacts rather than robust multimodal perception.
@@ 6.4 The Scaling Limitations in Multimodality
A persistent belief is that the architectural limitations of MLLMs can be overcome through scaling, i.e., adding more parameters, training on larger multimodal datasets, or increasing compute. This is justified for unimodal systems, where loss scales as a power law with model size, dataset size, and compute, leading to monotonic improvements (Kaplan et al., 2020; Hoffmann et al., 2022b). However, when vision and language are combined, scaling laws become fundamentally fragmented, introducing non-linearities that predictable unimodal scaling cannot accommodate.
@@ 6.4.1 Divergent Scaling Laws by Modality
$L_{\text{multi}}(N,D)=g(L_{\text{text}}(N,D),L_{\text{vision}}(N,D))$
@@
$\alpha_{\text{eff}}=\frac{\partial\log L_{\text{MLLM}}}{\partial\log N}\in\left[\min(\alpha_{\text{text}},\alpha_{\text{vision}}),\max(\alpha_{\text{text}},\alpha_{\text{vision}})\right]$
@@
The system exhibits anti-scaling: increased data in one modality can paradoxically worsen overall performance if modality imbalance is worsened. While modality balancing approaches, such as (Liu et al., 2025a), propose a solution during preference tuning on a much smaller dataset, it is unclear how such an approach can be applied for pretraining at scale.
@@ 6.4.3 Data Alignment Noise
A critical phenomenon in large-scale multimodal datasets is the non-linear compounding of the alignment noise. Let $p_{\text{misalign}}$ is the fraction of misaligned image-text pairs. The effective noise in the fused pair scales as:
$\epsilon_{\text{alignment}}\sim 1-\prod_{i=1}^{|D|}(1-p_{\text{misalign}})_{i}\approx 1-(1-p_{\text{misalign}})^{|D|}$
@@ 6.5.1 Promising Directions
Neuro-Symbolic Grounding. Neuro-symbolic approaches integrate multimodal embeddings with explicit symbolic reasoning to enforce compositional structure and causal constraints (Kamali et al., 2024). (Sehgal et al., 2024) introduced Cosmos framework, which uses neurosymbolic scene encodings, representing entities with neural vectors and composable symbolic attributes, combined via neurosymbolic attention that enforces interaction rules. Unlike end-to-end neural models, objects remain discrete with explicit attributes, and relations are encoded as symbolic rules rather than implicit attention patterns. Language-Regularized Concept Learners (LARC) (Feng et al., 2024a) leverage LLMs to impose linguistic constraints (word relations and semantic hierarchies) on neuro-symbolic 3D visual representations, making the predictions consistent with both visual evidence and symbolic rules. NeSyGPT (Cunnington et al., 2024) combines foundation model knowledge with neuro-symbolic reasoning by fine-tuning a vision-language foundation model to extract symbolic features, then learning answer set programs to solve downstream tasks, reducing labeling requirements while maintaining interpretability.
@@ 7 Evaluation Benchmark Limitations
Having established that LLM failures stem from five fundamental theoretical limits, namely hallucination, context compression, reasoning degradation, retrieval fragility, and multimodal misalignment, a natural question emerges - why do current benchmarks suggest continuous progress despite these intrinsic ceilings? This section demonstrates that contemporary evaluation practices systematically obscure these limits rather than measure them. Data contamination inflates scores by rewarding memorization over reasoning; judge bias incentivizes confident fabrication aligned with evaluator priors; compute-agnostic metrics hide the cost of marginal gains; and evaluation instability masks the saturation of genuine capability. Together, these artifacts create an illusion of progress that conflates benchmark score increases with fundamental capability advances. Understanding evaluation limitations is thus essential to interpreting where scaling helps versus where it merely exploits measurement artifacts. Despite substantial progress in evaluation frameworks, benchmarking LLMs remains fundamentally constrained by issues of (i) data contamination, (ii) judge and protocol bias, (iii) compute and efficiency, (iv) stability, and (v) equity and safety.
@@
Let $\mathcal{D}_{\text{train}}$ denote the pre-training dataset, and $\mathcal{B}$ the benchmark dataset used for evaluation. Let $\mathcal{D}^{\text{info}}_{\text{train}}$ and $\mathcal{B}^{\text{info}}$ represent the sets of all informational content (text, semantics, or paraphrased variants) contained in $\mathcal{D}_{\text{train}}$ and $\mathcal{B}$, respectively. Benchmark Data Contamination (BDC) occurs when there exists overlap between $\mathcal{D}_{\text{train}}$ and $\mathcal{B}$, causing the model to have prior exposure to evaluation data or semantically related knowledge. This overlap can substantially bias evaluation outcomes. (Xu et al., 2025a) quantifies the contamination risk as:
@@
An empirical study by (Lunardi et al., 2025) identified potential data contamination, especially in older benchmarks, indicating memorization rather than true understanding. This underscores the need for contamination-resistant benchmarks that are temporally updated and are semantically novel. The TS-Guessing protocol (Deng et al., 2024) and PaCoST statistical framework (Zhang et al., 2025b) enable contamination detection, though with inherent limitations in identifying sophisticated paraphrasing or semantic equivalence. Longitudinal analysis of contamination shows that performance drops after models’ training cutoff dates provide temporal evidence of contamination (Roberts et al., 2024). Beyond detection, recent work has explored mitigation strategies, with mixed results regarding the effectiveness of decontamination approaches (Sun et al., 2025). More fundamentally, (Chen et al., 2025b) introduced a dynamic data generation method to benchmark reasoning capabilities. Similarly, LiveBench (White et al., 2025) introduces a benchmark with automatically generated, frequently-updated questions and objective scoring, substantially reducing contamination risk through temporal dynamics.

### v2

@@ 6.3.2 Grounding and Reasoning Failures
Even when MLLMs integrate textual and other information (image, video, audio, etc.), the learned associations are predominantly spurious (Hosseini et al., 2025), which means that the associations are mere correlations in training data rather than causal or compositional structures. This manifests most clearly in spatial reasoning tasks, where models learn frequent co-occurrences (e.g., “cat on sofa”) but not the underlying geometric or physical relations (Rahmanzadehgervi et al., 2025; Hou et al., 2025). Let $\mathcal{S}=\{o_{1},\ldots,o_{n},r_{ij}\}$ denote a visual scene with objects $o_{i}$ and spatial relations $r_{ij}$ (e.g., “on top of,” “left of,” “inside”). An ideally grounded model would learn a representation $\phi(\mathcal{S})$ that factorizes as:
@@
where $\otimes$ denotes compositional combination and $\phi_{\text{rel}}$ explicitly encodes geometric constraints (e.g., vertical displacement for “on top of”). However, MLLM embeddings instead capture:
$\psi(\mathcal{S})\approx\sum_{i,j}w_{ij}\cdot\mathbbm{1}[\text{co-occur}(o_{i},o_{j})]$
where $w_{ij}$ are learned weights reflecting dataset co-occurrence frequencies and $\mathbbm{1}[\text{co-occur}(o_{i},o_{j})]$ indicates whether objects $o_{i}$ and $o_{j}$ appear together in training images.
@@
This spurious grounding is evident in compositional reasoning benchmarks. When presented with novel spatial configurations (e.g., “a sofa on top of a cat,” inverting the typical arrangement), MLLMs produce nonsensical outputs or refuse to generate descriptions, because the training data contains no instances of this configuration (Thrush et al., 2022; Yuksekgonul et al., 2023). The model learns a statistical pattern, not a geometric relationship that can be flexibly recombined. Research on compositional generalization shows that multimodal models often perform poorly when tested on compositional reasoning tasks that require understanding novel combinations of familiar concepts (Thrush et al., 2022). This limitation arises because the number of possible multi-way combinations grows exponentially, while training coverage is inherently limited (Lake et al., 2017; Bahdanau et al., 2019). It has been shown that even when models are explicitly trained using reinforcement learning (RL), compositional gaps with visual inputs remain (Li et al., 2025a). The failure is architectural: adding visual tokens simply expands the space of spurious correlations without imposing compositional structure.
@@
$\mathrm{Var}[Y\mid V,T]=\mathrm{Var}_{V}[Y\mid V]+\mathrm{Var}_{T}[Y\mid T]+2\,\mathrm{Cov}_{V,T}[Y].$
@@
Heavy reliance on synthetic captions or web-scraped image-text pairs causes MLLMs to overfit anthropomorphic and stylistic priors, propagating cognitive biases (Shumailov et al., 2023). Human-annotated captions often impose subjective interpretations (e.g., “the dog looks happy”), while synthetic captions generated by earlier models amplify errors across generations. As a result, MLLMs reproduce these biases, producing homogenized outputs that reflect training artifacts rather than robust multimodal perception.
@@ 6.4 The Scaling Limitations in Multimodality
A persistent belief is that the architectural limitations of MLLMs can be overcome through scaling, i.e., adding more parameters, training on larger multimodal datasets, or increasing compute. This may be justified for unimodal systems, where loss scales as a power law with model size, dataset size, and compute, leading to monotonic improvements (Kaplan et al., 2020; Hoffmann et al., 2022b). However, when vision and language are combined, scaling laws become fundamentally fragmented, introducing non-linearities that predictable unimodal scaling cannot accommodate.
@@ 6.4.1 Divergent Scaling Laws by Modality
$L_{\text{multi}}(N,D)=g(L_{\text{text}}(N,D),L_{\text{vision}}(N,D)).$
@@
$\alpha_{\text{eff}}=\frac{\partial\log L_{\text{MLLM}}}{\partial\log N}\in\left[\min(\alpha_{\text{text}},\alpha_{\text{vision}}),\max(\alpha_{\text{text}},\alpha_{\text{vision}})\right].$
@@
The system exhibits anti-scaling: increased data in one modality can paradoxically worsen overall performance if modality imbalance is worsened. While modality balancing approaches, such as Liu et al. (2025a), propose a solution during preference tuning on a much smaller dataset, it is unclear how such an approach can be applied for pretraining at scale.
@@ 6.4.3 Data Alignment Noise
A critical phenomenon in large-scale multimodal datasets is the non-linear compounding of the alignment noise. Let $p_{\text{misalign}}$ be the fraction of misaligned image-text pairs. The effective noise in the fused pair scales as:
$\epsilon_{\text{alignment}}\sim 1-\prod_{i=1}^{|D|}(1-p_{\text{misalign}})_{i}\approx 1-(1-p_{\text{misalign}})^{|D|}.$
@@ 6.5.1 Promising Directions
Neuro-Symbolic Grounding. Neuro-symbolic approaches integrate multimodal embeddings with explicit symbolic reasoning to enforce compositional structure and causal constraints (Kamali et al., 2024). Sehgal et al. (2024) introduced the Cosmos framework, which uses neurosymbolic scene encodings, representing entities with neural vectors and composable symbolic attributes, combined via neurosymbolic attention that enforces interaction rules. Unlike end-to-end neural models, objects remain discrete with explicit attributes, and relations are encoded as symbolic rules rather than implicit attention patterns. Language-Regularized Concept Learners (LARC) (Feng et al., 2024a) leverage LLMs to impose linguistic constraints (word relations and semantic hierarchies) on neuro-symbolic 3D visual representations, making the predictions consistent with both visual evidence and symbolic rules. NeSyGPT (Cunnington et al., 2024) combines foundation model knowledge with neuro-symbolic reasoning by fine-tuning a vision-language foundation model to extract symbolic features, then learning answer set programs to solve downstream tasks, reducing labeling requirements while maintaining interpretability.
@@ 7 Evaluation Benchmark Limitations
Having established that LLM failures stem from five fundamental theoretical limits, namely hallucination, context compression, reasoning degradation, retrieval fragility, and multimodal misalignment, a natural question emerges—why do current benchmarks suggest continuous progress despite these intrinsic ceilings? This section demonstrates that contemporary evaluation practices systematically obscure these limits rather than measure them. Data contamination inflates scores by rewarding memorization over reasoning; judge bias incentivizes confident fabrication aligned with evaluator priors; compute-agnostic metrics hide the cost of marginal gains; and evaluation instability masks the saturation of genuine capability. Together, these artifacts create an illusion of progress that conflates benchmark score increases with fundamental capability advances. Understanding evaluation limitations is thus essential to interpreting where scaling helps versus where it merely exploits measurement artifacts. Despite substantial progress in evaluation frameworks, benchmarking LLMs remains fundamentally constrained by issues of (i) data contamination, (ii) judge and protocol bias, (iii) compute and efficiency, (iv) stability, and (v) equity and safety.
@@
Let $\mathcal{D}_{\text{train}}$ denote the pre-training dataset, and $\mathcal{B}$ the benchmark dataset used for evaluation. Let $\mathcal{D}^{\text{info}}_{\text{train}}$ and $\mathcal{B}^{\text{info}}$ represent the sets of all informational content (text, semantics, or paraphrased variants) contained in $\mathcal{D}_{\text{train}}$ and $\mathcal{B}$, respectively. Benchmark Data Contamination (BDC) occurs when there exists overlap between $\mathcal{D}_{\text{train}}$ and $\mathcal{B}$, causing the model to have prior exposure to evaluation data or semantically related knowledge. This overlap can substantially bias evaluation outcomes. Xu et al. (2025a) quantifies the contamination risk as:
@@
An empirical study by Lunardi et al. (2025) identified potential data contamination, especially in older benchmarks, indicating memorization rather than true understanding. This underscores the need for contamination-resistant benchmarks that are temporally updated and are semantically novel. The TS-Guessing protocol (Deng et al., 2024) and PaCoST statistical framework (Zhang et al., 2025b) enable contamination detection, though with inherent limitations in identifying sophisticated paraphrasing or semantic equivalence. Longitudinal analysis of contamination shows that performance drops after models’ training cutoff dates provide temporal evidence of contamination (Roberts et al., 2024). Beyond detection, recent work has explored mitigation strategies, with mixed results regarding the effectiveness of decontamination approaches (Sun et al., 2025). More fundamentally, Chen et al. (2025b) introduced a dynamic data generation method to benchmark reasoning capabilities. Similarly, LiveBench (White et al., 2025) introduces a benchmark with automatically generated, frequently-updated questions and objective scoring, substantially reducing contamination risk through temporal dynamics.